OpenStudy (anonymous):

A historian took a count of the number of people in a gold rush town for six years in the 1870s. Find a quadratic function that models the data as a function of x, the number of years since 1870. Use the model to estimate the number of people in the town in 1878. P (x) = –6x2 + 25x + 475; 291 people P(x) = –6x2 + 25x + 475; 356 people P(x) = –2x2 – 25x + 425; 152 people P(x) = –2x2 – 25x + 425; 97 people

OpenStudy (anonymous):

Let me post the table.

OpenStudy (anonymous):

take your time and hurry up

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

choices?

OpenStudy (anonymous):

P (x) = –6x2 + 25x + 475; 291 people P(x) = –6x2 + 25x + 475; 356 people P(x) = –2x2 – 25x + 425; 152 people P(x) = –2x2 – 25x + 425; 97 people

OpenStudy (anonymous):

P(x) = –6x2 + 25x + 475; 356 people